OpenStudy (anonymous):

If a sample of HNO3 is found to contain 6.74 · 1024 N atoms how many moles of H atoms are present?

OpenStudy (anonymous):

Molar Mass (g/mol) HNO3 63.013 N atom 14.007 H atom 1.0080 O atom 15.999 Avogadro's No. (mol-1) 6.022×1023

ganeshie8 (ganeshie8):

In HNO3, number of H atoms = number of N atoms number of moles of H = \[\frac{number of\ H\ atoms}{Avagadro\ number}\] \[ =\frac{6.74 * 10^{24}}{6.022 * 10^{23}}\] = 11.2 moles